OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3) A.-2/3x+3 b.-2/3x+5/3 c.y=2x-21 d.y=-2/3x-9

OpenStudy (anonymous):

it is A

OpenStudy (anonymous):

two lines are said to be parallel only if their slopes are equal...... equation of a line in point slope form is...... (y-y1)=m*(x-x1)....... where m is the slope and(x1,y1) is the point thriugh which the line passes..

OpenStudy (anonymous):

so rearrange the equation 2x + 3y = 5 3y = -2x + 5 y = -2/3x +5/3 so the gradient/slope is -2/3 and so now we can sub it in the equation of line formula because we have the point (9,-3) as well. try sub those in (y-y1)=m*(x-x1)

OpenStudy (anonymous):

I really dont have a clue on how to solve it